The population of a city increases @ 4% p.a. There is an additional annual increase of 4% of the population due to the influx of job seekers, find the % increase in population after 2 years ?
Answer by JimboP1977(311)

Assuming the population increase for the first year is 4% and the increase for second year is an ADDITIONAL 4% then:
x = population number
After 1 year population = x*1.04
After 2 years population = x*1.04*1.04*1.04 = x*1.04^3 = x*1.124864
So after 2 years the population has increases by (1-1.124864)*100 percent or 12.48% to 2 dp.